OpenStudy (anonymous):

Solve for x: 3x-1/4 = 5/11

OpenStudy (anonymous):

x = 31/132 or .234848484 etc

OpenStudy (anonymous):

3x=5/11+1/4 ===> x=31/321

OpenStudy (mimi_x3):